OpenStudy (anonymous):

find the degree measures of the angle whose tangent is √3

OpenStudy (amistre64):

since tangent = y/x ; or opp/adj, we can set this up in a triangle and then "remember" which basic format we are looking at :)

OpenStudy (amistre64):

trig is 98% remembering stuff

OpenStudy (amistre64):

\begin{array}c &&&&.\\ &&&.&.\\ &&.&&.&\sqrt{3}\\ &.&a&&.\\ .&.&.&.&.\\ &&1\\ \end{array}

OpenStudy (amistre64):

this matches the 60-30-90 triangle; where a = 60 degrees